OpenStudy (anonymous):

Write a system of two equations using x and y in each for the following. The sum of two numbers is 170 and their difference is 26. Find the two numbers.

OpenStudy (anonymous):

Given: y+x=170 y-x=26 so solve for one variable. y=170-x and plug it into the second equation y-x=170 (170-x)-x=26 170-2x=26 196=2x (put all like terms on one side. remember what you do to one side you have to do to the other side. I just added +2x on both sides and minus 26 on both sides) x=98 (now divide) and now you solve for y if x=98 then... y+x=170 y+98=170 y=72